Determine the wavelength and colour of visible light that contains enough energy to eject an electron off a metal with a work function of 3eV? Take 1 J = 6.24 x 1018 eV and h = 6.63 x 10-34 J.s
Answer #1

We know that

Energy E =hc/λ

c is the speed of light = 3.00×108m/s

λ is the wavelength

h is the Planck Constant which = 6.63×10−34 J.s

λ = hc / E =  6.63×10−34 J.s * 3.00×108m/s / (3 * 1.6*10-19) = 414 nm

The color is violet since the range of violet is 390 nm to 430 nm